OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). Answer Choices: y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

@Hamoody1996 hi, could you help me with this question? :]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i mean the third answer ;)

OpenStudy (anonymous):

y=3x-14

OpenStudy (anonymous):

oh :3 could ya explain how you got the answer please?

OpenStudy (anonymous):

sure since they are parallel so they have the same slope (m) and the slope equals the first derivative for the equation ( 3x - y = -3) so \[m=\frac{ dy }{ dx } = 3\] and by applying the law of the slope (\[m=\frac{ y-y1 }{ x-x1 }\] and we apply the point (4 , -2) to get the equation feel free to ask any other question :)

OpenStudy (anonymous):

thank you! <3 Select the equation of a line that is parallel to the line on the graph and passes through the point (2, 3). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = - 1 over 2 x + 4

OpenStudy (anonymous):

@Hamoody1996

OpenStudy (anonymous):

first you have to find the equation of the line which is \[y=-1-\frac{ x }{ 2 }\] and by finding the derivative to and do all the previous steps...